# Emerging Platforms Meet Emerging LLMs: A Year-Long Journey of Top-Down Development

### v2

@@ III-B Prerequisites
While TapML is general, it makes the following lightweight assumptions to the ML framework using TapML:
@@
Compute graph abstraction: TapML requires the developing system to model ML computation using a compute graph, i.e., a directed graph of tensor operators, each of which transforms a set of input tensors to output tensors. The abstraction of the compute graph allows for straightforward instrumentation of input and output tensors during test carving (in Section III-C). Viewing an operator as a minimal computation unit also offers a decent granularity for implementation, validation, and offloading to the intended platform (in Section III-D). This assumption is reasonable as the concept of compute graphs has long existed in various major ML frameworks and compilers, such as TensorFlow [23] and PyTorch [5].
@@
III-C Test Carving
Figure 5: Carving intermediate inputs and outputs to a test case.
Test Carving [14] is a technique to automatically extract fine-grained unit tests from a given system-level test. Such unit tests are used to replay the intermediate behaviors (e.g., individual function invocations) that happened during the system test. To achieve this, carving techniques instrument [24] a system test to record invocations of desired functions along with their inputs and outputs. By knowing the function, inputs (e.g., arguments and context), and outputs (i.e., test oracle of differential testing [15]), a test carver synthesizes a unit test to reproduce the intermediate step of the system test.
In model deployment, a system test refers to the end-to-end validation of a full model, where the model is given some semantically meaningful inputs (e.g., pictures and sentences) and is expected to produce reasonable end-to-end outputs (e.g., correctly completing a motto). From the high level, TapML carves operator-level unit tests by instrumenting the end-to-end model execution on the source platform, to obtain reliable test inputs and reference outputs. Specifically, Figure 5 illustrates the carving process. First, the original compute graph on the source platform is modified such that each operator $\mathcal{F}_{S}$ is wrapped by an instrumentor, highlighted in the blue-green box. The instrumentation includes two instructions: (i) a pre-operator instruction to collect the operator call $\mathcal{F}$ and runtime inputs $\mathcal{I}$ feed to the instrumented operator; and (ii) a post-operator instruction to document the produced output $\mathcal{O}=\mathcal{F}_{S}(\mathcal{I})$ as the oracle from the source platform. Hence, each operator per execution forms a unit test as $t=(\mathcal{F},\mathcal{I},\mathcal{O})$ which can be replayed to validate the target-platform implementation. Such a source-to-target cross-checking mechanism also indirectly implements Differential Testing [15].
Notably, the model-wise inputs for bootstrapping the instrumented system test should be semantically meaningful, e.g., real-world data from the evaluation test set. This is important as it allows developers to focus the precision regression over realistic inputs, whereas synthetic test inputs used in the bottom-up pipeline is inadequate to validate the numeric precision stability due to the inconsistent distribution to natural inputs, though convenient as it is challenging to specify meaningful high-dimensional embeddings as intermediate inputs.
III-D Gradual Target Offloading
@@ 25             $\langle\mathcal{I},\mathcal{O}\rangle\leftarrow t_{i}$
TapML employs the Gradual Target Offloading mechanism to progressively transfer the compute graph from a standalone host to a standalone target platform. Algorithm 1 details steps of the Gradual Target Offloading procedure. In the high level, Gradual Target Offloading takes a source-platform compute graph $\mathcal{M}_{S}$ as input and migrates it to a target-platform compute graph $\mathcal{M}_{T}$. Recall Section III-B that $\mathcal{M}_{S}$ is a reference implementation from a mature source platform, assumed to be available beforehand. Both $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$ can be initially represented using a general platform-agnostic representation shown in Figure 4 but they are specialized to different code targets. Specifically, $\mathcal{M}_{S}$ can be defined by a list of operators $\mathbb{F}_{S}$ and their topology information $\mathcal{G}$ (Algorithm 1). Therefore, to obtain $\mathcal{M}_{T}$ the goal is to transform $\mathbb{F}_{S}$ to $\mathbb{F}_{T}$ starting from an empty set (Algorithm 1). To start with, we perform test carving (Section III-C) to get a test suite where each operator maps to its corresponding unit tests (Algorithm 1). For each operator $\mathcal{F}_{i}$ in $\mathbb{F}_{S}$ we run following migration steps: (i) We draft an initial implementation for $\mathcal{F}_{i}$ on the target platform, i.e., $\mathcal{F}_{Ti}$ (Algorithm 1); (ii) we test and debug the implementation of $\mathcal{F}_{Ti}$ using the carved tests until the operator-wise validation passes (Algorithms 1 and 1); (iii) after operator validation, we perform model-wise validation over a hybrid compute graph $\langle\mathbb{F}^{*},\mathcal{G}\rangle$, by substituting corresponding parts in $\mathbb{F}_{S}$ with implemented target-platform operators, i.e., $\left\{\mathbb{F}_{T}\cup\{\mathcal{F}_{Ti}\}\right\}$ (Algorithms 1, 1 and 1); and (iv) lastly we append the tested $\mathcal{F}_{Ti}$ to $\mathcal{M}_{T}$ to finalize the migration of $\mathcal{F}_{i}$ (Algorithm 1). As such, once all target-platform operators are implemented and tested (i.e., $\mathbb{F}_{T}$), we wrap up the procedure by returning a compute graph $\mathcal{M}_{T}$ fully supported on the target platform (Algorithms 1 and 1).
Figure 6: Sample validation code using TapML APIs.
III-E Universal Runtime
@@
As is shown in Figure 4, we need a runtime system to transfer data (e.g., device code and test inputs) and perform differential testing. To this end, we build a universal runtime to unify these procedures to facilitate cross-platform data transfer and operation. We argue that it is crucial to unify related behaviors through a standardized interface, which provides a reliable programming model and maximizes code reuses.
@@
Exemplification. Figure 6 demonstrates the interface via an example of differential testing on CUDA and WebGPU implementations. Specifically, the tester takes a transformer_block operator module and a set of computational inputs as the function input (Line 3), based on which we compile the module to both CUDA and WebGPU platforms and perform differential testing (Line 21). In Line 4, the local function run compiles the transformer_block module on a given device, returning the results of running the compiled module over the given inputs. More specifically, we first compile the target device code locally (Line 6) and upload the compiled device code to the remote device via tapml.upload (Line 7), which returns a handle (i.e., reference) to the remote library. Next, we convert and transfer the test inputs from the local host to the device platform as tml_inputs in Line 10. To obtain the output on the device, we also need to allocate space for the output tensor tml_output in Line 12. As such, in Line 15 and 16 we call the remote device function which transforms tml_inputs to tml_output, which are transferred back to the local host in Line 19 to perform concrete testing.
@@
Implementation. The core of the runtime is communication, i.e., how to transfer data and instructions across multiple platforms. Consequently, we develop a lightweight Remote Procedure Call (RPC) framework to manage host-device data transfer and perform device function calling. Additionally, error handling is a critical yet challenging aspect of universal runtime systems. When an error occurs, the runtime is designed to collect error information as much as possible, with complete trackback information, and transmit these debugging messages back to the host before device termination.
IV Evaluation
@@
  • RQ1 (Section IV-B): How is the overall model support and timeline by using TapML?
  • RQ2 (Section IV-C): How is the performance and overhead of using the universal runtime component?
  • RQ3 (Section IV-D): What are the bugs and challenges when developing emerging ML systems?
  • RQ4 (Section IV-E): What are the lessons and best practices of using TapML?
IV-A Setup
@@
row2,4,5,8,9=bg=azure9, colspec=Q[c,white] — Q[c] Q[c] Platform Hardware
@@
WebGPU Apple Silicon M1 Pro
(MacBook Pro)
@@
(Xiaomi 14 Pro)
@@
(iPhone 14 Pro)
ROCm Radeon 7900 XTX
TABLE I: Mature and emerging platforms in TapML.
Platforms. To demonstrate the versatility and adaptability of TapML across a diverse range of configurations, we evaluate TapML over platforms shown in Section IV-A. The corresponding hardware selection spans from a GPU-intensive workstation to energy-efficient mobile devices. Furthermore, we consider a rich set of mature and emerging runtime platforms below:
  1. 1.
    CUDA [25]: A parallel computing platform and application programming interface model for NVIDIA GPUs.
  2. 2.
    ROCm [26]: An open-source platform by AMD for GPU-accelerated computing, analogous to NVIDIA’s CUDA but for AMD GPUs. It provides a comprehensive suite of tools and libraries for high-performance, parallel computing applications across various domains.
  3. 3.
    Metal [11]: Apple’s low-level API for optimizing graphics and computing on Apple ecosystems such as iOS and macOS, offering high performance and efficiency.
  4. 4.
    WebGPU [12]: A web standard providing cross-platform, low-level graphics and computation APIs and capabilities on web browsers.
  5. 5.
    Vulkan [27]: A cross-platform graphics and compute API, providing high-efficiency, low-overhead access to modern GPUs on various devices.
  6. 6.
    OpenCL [28]: An open, royalty-free standard for cross-platform, parallel programming of diverse processors found in personal computers, servers, mobile devices, and embedded platforms.
Base framework. We experiment TapML based on Apache TVM [6], an open-source ML compiler stack with robust support for CPU and NVIDIA GPU platforms. Specifically, we build a universal ML inference engine targeting emerging models and platforms as part of the MLC-LLM project at https://github.com/mlc-ai/mlc-llm on top of TVM.
IV-B RQ1: Support Overview
@@
StarCoder-1B Apr 24 2023 June 23 2023 430.2 157.8 388.9 86.0 49.1 23.1 38.1
RedPajama-3B May 5 2023 May 19 2023 283.0 131.2 245.1 47.5 29.7 14.1 19.5
InternLM-7B July 6 2023 Feb 28 2024 138.8 96.3 148.4 26.0 15.7 8.3 3.2
WizardMath-7B Aug 11 2023 Aug 12 2023 147.2 106.4 168.6 27.4 17.2 8.8 5.4
Baichuan2-7B Sept 6 2023 Feb 16 2024 129.1 97.1 153.5 26.9 14.1 10.8 1.9
TinyLlama-1.1B Sept 16 2023 Sept 21 2023 577.9 279.2 487.5 103.3 63.1 29.9 42.3
Mistral-7B Sept 27 2023 Sept 28 2023 128.1 85.1 139.5 24.7 15.5 6.3 4.1
StableLM-3B Sept 29 2023 Oct 3 2023 281.2 135.8 240.8 47.4 34.2 19.9 19.6
@@
In a year-long effort (starting from April 2023), we based on TapML have deployed 82 models (i.e., validated) of 17 model architectures over 5 emerging platforms. To our knowledge, to date TapML offers the most comprehensive range of platform support. More importantly, TapML also becomes the fastest and even sole supporter on a line of important scenarios:
  1. 1.
@@
    The first LLM on mobile GPUs: Thanks to TapML, we are the first in open source to successfully deploy LLMs on mobile GPUs of Android and iOS devices. By maximizing the efficiency of mobile GPUs via Metal and OpenCL, running large LLMs on mobile phones becomes a possible and decent experience, e.g., running Baichuan2-7B on Snapdragon 8 Gen 3 at over 10 tokens per second. Our pioneering work motivates the community to develop native and mobile LLM applications without the fear of sacrificing user privacy.
  2. 2.
@@
    The first LLM on WebGPU: TapML allows us to be the first in public to run LLMs on WebGPU, despite its status as an experimental feature in Chrome at the time of development. Enabling emerging ML on WebGPU is of significant broader impact as it allows everyone to run powerful models as easily as opening a web page. Notably, nearly one year after our effort, Google introduced the MediaPipe [29] via LLM Inference API on TensorFlow.js [16] for WebGPU and mobile platforms. We believe TapML approach can complement and speed up the development of these related efforts.
  3. 3.
@@
    The only Orange Pi LLM: To date, we are the only team that has successfully deployed LLMs on the Orange Pi 5 [30], an edge device and embedded system in \$100. Yet, we achieve a practical operational speed of 2.5 tokens per second for running Llama2-7B and 5 tokens per second for RedPajama-3B using the OpenCL APIs to maximize the edge GPU on Orange Pi. This inspires the community to develop affordable LLM applications on edge scenarios.
@@
Thanks to the techniques and design choices in TapML, these highlights are achieved by our small team. Through test carving TapML frees developers from writing tedious test cases. Meanwhile, the gradual offloading mechanism in TapML minimizes the mental burden of developers by allowing them to focus on the minimal implementation unit of single operators each time until it is fully validated. Lastly, the universal runtime becomes a huge time saver since interacting with such emerging and under-developed platforms is made as easy as calling Python functions locally.
More specifically, LABEL:tab:model_support details the availability and performance of a selection of models and platform supports. Thanks to TapML, we quickly deploy recent and popular models within a short period, e.g., up-to same-day support, demonstrating the effectiveness of TapML in optimizing development productivity. More importantly, all of the deployed models are extensively validated by carved tests from TapML as well as common end-to-end tests and achieve outstanding processing performance to serve these models efficiently.
IV-C RQ2: Overhead of TapML
@@
row2,4,6,8=bg=azure9, colspec=Q[c,white] — Q[l] Q[r] Q[c] Q[r] Development step Platform Cost O(1)?

### v3

@@ 3.2. Prerequisites
While TapML is general, it makes the following lightweight assumptions to the deployment process:
@@
Compute graph abstraction: TapML requires the developing system to model ML computation using a compute graph, i.e., a directed graph of tensor operators, each of which transforms a set of input tensors to output tensors. The abstraction of the compute graph allows for straightforward instrumentation of input and output tensors during test carving (in Section 3.3). Viewing an operator as a minimal computation unit also offers a decent granularity for implementation, validation, and offloading to the intended platform (in Section 3.4). This assumption is reasonable as the concept of compute graphs has long existed in various major ML frameworks and compilers, such as TensorFlow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019).
@@
3.3. Test Carving
Figure 4. Carving intermediate inputs and outputs to a test case.
Test Carving (Elbaum et al., 2006) is a technique to automatically extract fine-grained unit tests from a given system-level test. The extracted unit tests are used to replay the intermediate behaviors (e.g., individual function calls) that happened during the system test. To achieve this, test carving instruments (Huang, 1978) a system test by recording invocations of desired functions and their corresponding inputs and outputs. By knowing the function, inputs (e.g., arguments and context), and outputs (i.e., test oracle of differential testing (McKeeman, 1998)), a test carver synthesizes a unit test to reproduce the expected intermediate step of the system test.
In model deployment, a system test refers to the end-to-end validation of a full model, where the model is given some semantically meaningful inputs (e.g., pictures and sentences) and is expected to produce reasonable end-to-end outputs (e.g., correctly completing a motto). From the high level, TapML carves operator-level unit tests by instrumenting the end-to-end model execution on the source platform, to obtain reliable test inputs and reference outputs. Specifically, Figure 4 illustrates the carving process. First, the original compute graph on the source platform is modified such that each operator $\mathcal{F}_{S}$ is wrapped by an instrumentor, highlighted in the blue-green box. The instrumentation includes two instructions: (i) a pre-operator instruction to collect the operator call $\mathcal{F}$ and runtime inputs $\mathcal{I}$ feed to the instrumented operator; and (ii) a post-operator instruction to document the produced output $\mathcal{O}=\mathcal{F}_{S}(\mathcal{I})$ as the oracle from the source platform. Hence, each operator per execution forms a unit test as $t=(\mathcal{F},\mathcal{I},\mathcal{O})$ which can be replayed to validate the target-platform implementation. Such a source-to-target cross-checking mechanism also indirectly implements Differential Testing (McKeeman, 1998).
Notably, the model-wise inputs for bootstrapping the instrumented system test should be semantically meaningful, e.g., real-world data from the evaluation test set. This is important as it allows developers to focus the precision regression over realistic inputs, whereas synthetic test inputs, though meaningful for defending adversarial attacks (Pei et al., 2017), are inadequate to validate the numeric precision stability due to the inconsistent distribution of natural inputs, though convenient as it is challenging to specify meaningful high-dimensional embeddings as intermediate inputs.
3.4. Gradual Target Offloading
@@ 25             $\langle\mathcal{I},\mathcal{O}\rangle\leftarrow t_{i}$
TapML employs the Gradual Target Offloading mechanism to progressively transfer the compute graph from a standalone host to a standalone target platform. Algorithm 1 details steps of the Gradual Target Offloading procedure. In the high level, Gradual Target Offloading takes a source-platform compute graph $\mathcal{M}_{S}$ as input and migrates it to a target-platform compute graph $\mathcal{M}_{T}$. Recall Section 3.2 that $\mathcal{M}_{S}$ is a reference implementation from a mature source platform, assumed to be available beforehand. Both $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$ can be initially represented using a general platform-agnostic representation shown in Figure 3, but they are specialized to different code targets. Specifically, $\mathcal{M}_{S}$ can be defined by a list of operators $\mathbb{F}_{S}$ and their topology information $\mathcal{G}$ (Algorithm 1). Therefore, to obtain $\mathcal{M}_{T}$ the goal is to transform $\mathbb{F}_{S}$ to $\mathbb{F}_{T}$ starting from an empty set (Algorithm 1). To start with, we perform test carving (Section 3.3) to get a test suite where each operator maps to its corresponding unit tests (Algorithm 1). For each operator $\mathcal{F}_{i}$ in $\mathbb{F}_{S}$ we run following migration steps: (i) We draft an initial implementation for $\mathcal{F}_{i}$ on the target platform, i.e., $\mathcal{F}_{Ti}$ (Algorithm 1); (ii) we test and debug the implementation of $\mathcal{F}_{Ti}$ using the carved tests until the operator-wise validation passes (Algorithms 1 and 1); (iii) after operator validation, we perform model-wise validation over a hybrid compute graph $\langle\mathbb{F}^{*},\mathcal{G}\rangle$, by substituting corresponding parts in $\mathbb{F}_{S}$ with implemented target-platform operators, i.e., $\left\{\mathbb{F}_{T}\cup\{\mathcal{F}_{Ti}\}\right\}$ (Algorithms 1, 1 and 1); and (iv) lastly we append the tested $\mathcal{F}_{Ti}$ to $\mathcal{M}_{T}$ to finalize the migration of $\mathcal{F}_{i}$ (Algorithm 1). As such, once all target-platform operators are implemented and tested (i.e., $\mathbb{F}_{T}$), we wrap up the procedure by returning a compute graph $\mathcal{M}_{T}$ fully supported on the target platform (Algorithms 1 and 1).
Figure 5. Sample validation code using TapML APIs.
3.5. Universal Runtime
@@
As is shown in Figure 3, we need a runtime system to transfer data (e.g., device code and test inputs) and perform differential testing. To this end, we build a universal runtime to unify these procedures to facilitate cross-platform data transfer and operation. We argue that it is crucial to unify related behaviors through a standardized interface, which provides a reliable programming model and maximizes code reuse.
@@
Exemplification. Figure 5 demonstrates the interface via an example of differential testing on CUDA and WebGPU implementations. Specifically, the tester takes a transformer_block operator module and a set of computational inputs as the function input (Line 3), based on which we compile the module to both CUDA and WebGPU platforms and perform differential testing (Line 21). In Line 4, the local function run compiles the transformer_block module on a given device, returning the results of running the compiled module over the given inputs. More specifically, we first compile the target device code locally (Line 6) and upload the compiled device code to the remote device via tapml.upload (Line 7), which returns a handle (i.e., reference) to the remote library. Next, we convert and transfer the test inputs from the local host to the device platform as tml_inputs in Line 10. To obtain the output on the device, we also need to allocate space for the output tensor tml_output in Line 12. As such, in Line 15 and 16 we call the remote device function, which transforms tml_inputs to tml_output, which are transferred back to the local host in Line 19 to perform concrete testing.
@@
Implementation. The core of the runtime is communication, i.e., how to transfer data and instructions across multiple platforms. Consequently, we developed a lightweight Remote Procedure Call (RPC) framework to manage host-device data transfer and perform device function calling. Additionally, error handling is a critical yet challenging aspect of universal runtime systems. When an error occurs, the runtime is designed to collect error information as much as possible, with complete trackback information, and transmit these debugging messages back to the host before device termination.
4. Evaluation
@@
  • RQ1 (Section 4.2): How is the overall model support progress by using TapML?
  • RQ2 (Section 4.3): How TapML improves developer productivity in model debugging?
  • RQ3 (Section 4.4): What is the overhead of applying TapML to model deployments?
  • RQ4 (Section 4.5): What do the bugs look like when deploying models on emerging platforms? How does the TapML approach help accelerate the troubleshooting process?
  • RQ5 (Section 4.6): What are the empirical lessons and best practices of using TapML?
4.1. Setup
Framework. TapML is built as the testing and debugging methodology and tooling behind the MLC-LLM (MLC team, 2025) project, which focuses on native deployment of LLM models on a diverse set of environments across cloud, edge, and emerging platforms like web browsers. Specifically, as MLC-LLM is built on top of Apache TVM (Chen et al., 2018), an open-source ML compiler infrastructure, our implementation of TapML relies on the TVM framework to perform various operations, such as manipulating models via its compiler intermediate representation (Lai et al., 2025).
@@ 4.1. Setup
row2,4,6=bg=azure9, colspec=Q[c,white] — Q[c] Q[c] Q[c] Platform Hardware RAM Size
@@
WebGPU Apple M1 Pro (Macbook Pro) 16 GB
OpenCL Snapdragon 8 Gen 3 (Xiaomi 14 Pro) 16 GB
ROCm Radeon 7900 XTX 24 GB
@@
Apple A18 Pro (iPhone 16 Pro) 8 GB
Table 1. Platforms covered in our experimental setup.
Platforms. To demonstrate the versatility and adaptability of TapML across a diverse range of configurations, we evaluate TapML over platforms shown in Section 4.1. The corresponding hardware selection spans from a GPU-intensive workstation to energy-efficient mobile devices. Furthermore, we consider a rich set of mature and emerging runtime platforms below:
  1. (1)
    CUDA (Luebke, 2008): A parallel computing platform and application programming interface model for NVIDIA GPUs.
  2. (2)
    ROCm (AMD, 2024): An open-source platform by AMD for GPU-accelerated computing, analogous to NVIDIA’s CUDA but for AMD GPUs. It provides a comprehensive suite of tools and libraries for high-performance, parallel computing applications across various domains.
  3. (3)
    Metal (Apple, 2024): Apple’s low-level API for optimizing graphics and parallel computing on Apple ecosystems. Note that the development setups for iOS-based Metal and macOS-based Metal are quite different; yet, we count the two development pipelines as one platform.
  4. (4)
    WebGPU (Consortium, 2024): A web standard providing cross-platform, low-level graphics and computation APIs and capabilities on web browsers.
  5. (5)
    Vulkan (Sellers and Kessenich, 2016): A cross-platform graphics and compute API, providing high-efficiency, low-overhead access to modern GPUs on various devices.
  6. (6)
    OpenCL (Munshi, 2009): An open and royalty-free standard for cross-platform parallel programming of diverse processors found in servers, mobile devices, embedded platforms, etc.
4.2. RQ1: Platform and Model Support
{tblr}
@@ 4.2. RQ1: Platform and Model Support
Ours Others
WebGPU May 1, 2023 Mar 7, 2024 (Google MediaPipe)
OpenCL on Android May 1, 2023 Mar 7, 2024 (Google MediaPipe)
Metal on iOS May 1, 2023 Nov 30 2023 (Apple MLX)
Vulkan May 1, 2023 Sep 18, 2023 (GPT4All)
ROCm May 1, 2023 Aug 25, 2023 (llama.cpp)
Metal on macOS May 1, 2023 Jun 5, 2023 (llama.cpp)
Table 2. Timeline of emerging platform support availability across frameworks. Our support is up to 10 months earlier than other open-source and industrial teams.
Overview. Starting in April 2023, based on TapML, we have deployed 105 models of 27 model architectures over 5 emerging platforms, empowering emerging AI applications on mobile phones, browsers, servers, and personal laptops. Meanwhile, to our knowledge, Section 4.2 shows that we are the first in the open-source community to support such a wide range of emerging platforms. For example, we supported the mobile and browser platforms 10 months ahead of the MediaPipe engine from Google. Furthermore, LABEL:tab:modelsup demonstrates our productivity by showing the gap between the model release date and our full-platform support date. Our team typically integrates support for new model architectures within 30 days of release. Longer implementation times generally reflect strategic prioritization decisions due to our small team size rather than technical limitations. It is also worth noting that when we say “support”, it means the model can not only run on our target platforms but also pass our model-wise testing.
{talltblr}
@@
[…]
Mistral Sep 27, 2023 Sep 28, 2023 1
Orion Jan 20, 2024 Mar 4, 2024 44
[…]
Qwen2 Feb 5, 2024 Feb 13, 2024 8
StarCoder2 Feb 28, 2024 Jul 14, 2024 137
Qwen2Moe Mar 14, 2024 Apr 5, 2024 22
Phi3 Apr 23, 2024 Apr 25, 2024 2
Cohere May 23, 2024 Jun 25, 2024 33
Gemma2 Jun 27, 2024 Jul 20, 2024 23
InternLM2 Jul 3, 2024 Jul 7, 2024 4
[…]
Phi3 V May 19, 2024 Jul 15, 2024 57
@@
Mistral Mistral-7B-Instruct 141.2 83.0 156.1 24.1 19.5 10.6 11.6
Orion Orion-14B-Chat 74.0 41.9 85.7 13.9 OOM OOM OOM
MiniCPM MiniCPM-2B-128k 234.2 124.2 208.0 33.1 17.1 13.4 27.0
StarCoder2 StarCoder2-3B 234.5 106.6 223.2 36.7 28.5 12.9 18.6
Qwen2Moe Qwen1.5-MoE-A2.7B-Chat 246.1 147.4 257.0 46.9 17.3 12.4 OOM
Phi3 Phi-3.5-mini-instruct 226.7 142.2 239.8 32.8 32.1 11.1 16.4
Gemma2 Gemma-2-2B-it 279.4 139.7 247.7 39.4 21.4 12.7 28.6
InternLM2 InternLM2.5-1.8B-Chat 410.0 178.0 369.0 67.0 46.1 20.4 45.1
@@
Phi3 V Phi-3.5-vision-instruct 234.6 154.9 240.6 29.7 31.3 11.3 17.0
[…]
Llama-3.2-1B-Instruct (Draft)
Highlights. We exemplify a few of our support highlights below:
  1. (1)
@@
    The first LLM on mobile GPUs: Thanks to TapML, we are the first in open source to successfully deploy LLMs on mobile GPUs of Android and iOS devices. By maximizing the efficiency of mobile GPUs via Metal and OpenCL, running large LLMs on mobile phones becomes a possible and decent experience, e.g., running Llama3-8B on Snapdragon 8 Gen 3 at over 10 tokens per second. Our pioneering work motivates the community to develop native and mobile LLM applications without the fear of sacrificing user privacy.
  2. (2)
@@
    The first LLM on WebGPU: TapML allows us to be the first in public to run LLMs on WebGPU, despite its status as an experimental feature in Chrome at the time of development. Enabling emerging ML on WebGPU is of significant broader impact as it allows everyone to run powerful models as easily as opening a web page. Notably, 10 months after our effort, Google introduced the MediaPipe (Lugaresi et al., 2019) via LLM Inference API on TensorFlow.js (Google, 2024) for WebGPU and mobile platforms. We believe TapML approach can complement and speed up the development of these related efforts.
  3. (3)
@@
    The only LLM support on Orange Pi GPU: To date, we are the only team that has successfully deployed LLMs on the Orange Pi 5 (Orange Pi, 2024), an edge device and embedded system in \$100. Yet, we achieve a practical operational speed of 2.5 tokens per second for running Llama2-7B and 5 tokens per second for RedPajama-3B using the OpenCL APIs to maximize the edge GPU on Orange Pi. This inspires the community to develop affordable LLM applications on edge scenarios.
@@
Thanks to the productivity gain brought by TapML, we are able to achieve these highlights and support emerging platforms much earlier than industrial and other open-source teams, despite that we are a team with less than 10 core developers. Through test carving TapML frees developers from writing tedious test cases and ensures the test quality. Meanwhile, the gradual offloading mechanism in TapML minimizes the mental burden of developers by allowing them to focus on the minimal implementation unit of single operators each time until it is fully validated. Lastly, the universal runtime becomes a huge time saver since interacting with such emerging and under-developed platforms is made as easy as calling Python functions locally.
Deployment quality. We demonstrate the quality of our model deployments by showcasing their inference efficiency and benchmark accuracy. LABEL:tab:modelperf lists the inference performance of recent models we deployed and quantized in 4 bits. This result shows that our swift model supports also come with decent quality by being able to serve these models efficiently, even under resource-constrained scenarios. Beyond efficiency, our deployments have undergone rigorous and productive testing with TapML. Section 4.2 evaluates several recent models deployed by our and other CUDA-focused frameworks using code generation benchmarks, in which these models excel. Specifically, we evaluate them using HumanEval+ (Liu et al., 2024), a program synthesis benchmark to rigorously test if LLMs can generate functionally correct code. The results show that our deployed models perform comparably, trailing by no more than three tasks relative to the median number of passing tasks achieved by other frameworks’ deployments.
Notably, in the LLM domain, slight result discrepancy can be expected due to the nature of floating-point arithmetic — the exact result depends on the floating-point accuracy (e.g., 16-bit or even lower precisions are commonly used for LLMs) and the detailed computation orders (e.g., $a+(b+c)\neq(a+b)+c$) of the implementation. During decoding, minor differences in decoding probability between two output logits can lead to different, but likely valid, tokens being chosen. Furthermore, due to the casual nature of LLMs, one inconsistent token in the context can lead to entirely different subsequent sentences.
@@
[…]
CUDA ROCm Vulkan Metal
Qwen2.5-Coder-1.5B-Instruct 101 m107 109 110 110 109 109
Llama-3.2-3B-Instruct 92 98 m94 92 92 91 94
Gemma-2-2B-it 57 m58 61 65 67 65 67
Qwen2.5-0.5B-Instruct 42 54 m53 52 52 51 51
Table 3. Number of passing HumanEval+ tasks (164 in total) for recent models deployed by our framework and other CUDA-focused frameworks. Our deployments can either outperform (i.e., Gemma-2-2B-it) or match the medium of reference results (highlighted as $\square$”) within 3-task discrepancy; yet, the popular HuggingFace transformer library can present a notable discrepancy of nearly 10 tasks (i.e., Qwen2.5-0.5B-Instruct).
4.3. RQ2: Productivity Improvements
To evaluate how much TapML can improve developer productivity when debugging model deployments, we designed a controlled experiment using human study to compare TapML and bottom-up methods. In particular, our human study focuses on localizing buggy operators instead of fixing them, because the efficiency of fixing depends on specific expertise and is orthogonal to the deployment methods.
Evaluation setup. We recruited three volunteers to debug 10 cases, each of which includes (i) a real-world model including weights and operator information, (ii) a golden CUDA library to run the model, and (iii) a buggy Vulkan library with manually injected bugs in the low-level model IR. All of the three participants are Ph.D. students in computer science with over 3 to 6 years of experience in ML system engineering. Each participant is randomly assigned to 3 to 4 cases. For each case, they are asked to localize the list of problematic operators and we record the debugging time and accuracy of debugging outcomes. Specifically, for bottom-up debugging, volunteers construct operator-wise test cases to perform differential testing over the two libraries. They are allowed to access tools that are available in real-world development scenarios, including but not limited to using search engines and LLMs, and can create reusable scripts and tools to (partially) automate the workflow. For top-down debugging, participants can use the TapML tool set which can automatically perform gradual target offloading (Section 3.4) over all operators.
Result overview. Section 4.3 demonstrates that TapML brings improvements in not only productivity but also quality in debugging. Despite the bottom-up baseline on average costing 31 minutes per model to implement tests to identify bugs, it incurs a false positive rate as high as 58.2% and a false negative rate of 5.5%. In contrast, TapML can quickly localize bugs in one minute with no false negatives and a light false positive rate of 10.3%. It is also worth noting that case #1 took significantly longer than the others because it is the first-time setup from the volunteer with the least experience in the toolchain.
False positives. The bottom-up approach struggled with a high number of false positives (32 in total across all models), with volunteers frequently reporting non-buggy operators. These false positives are frequent when doing bottom-up debugging for two reasons: (i) models are often quantized for efficient deployment and thus sensitive to the numerical distribution of inputs; and (ii) in bottom-up testing, developers often use random and unrealistic out-of-domain inputs that are not considered by the quantization. In contrast, the test inputs used in TapML are derived from real-world user inputs to meet the input expectation of model quantization, resulting in only 3 false positives from merely 1 model (i.e., case #6). Nonetheless, TapML is not perfect as it also yields 3 false positives in case #6, which uses a complex Mixture-of-Experts (MoE) model.
False negatives. The model behind case #10 is an LLM whose Vulkan implementation of paged attention (Kwon et al., 2023) is buggy. These bugs are missed in bottom-up debugging as our volunteers failed to implement tests for the paged attention kernel. Implementing paged attention is challenging as it requires a real-world historical key-value (KV) cache with complex page management. Using random inputs for this kernel usually leads to incompatible configurations and data, crashing the kernel due to API misuse. On the other hand, TapML uses realistic KV caches and configurations derived from the golden library implementation, leading to zero observed false negatives.
Overall, these findings confirm that TapML’s top-down design substantially improves the speed and quality of bug localization, providing developers with precise and actionable information about problematic operators, eliminating the need for extensive manual checking and minimizing the debugging scope.
{tblr}
[…]
ID # Unique Op. # Bugs Min. # Reports # FP # FN Min. # Reports # FP # FN
1 71 1 74 6 5 0 1 1 0 0
2 71 4 22 5 1 0 1 4 0 0
3 71 11 13 11 0 0 1 11 0 0
4 65 2 29 5 3 0 1 2 0 0
5 67 2 21 10 8 0 1 2 0 0
6 96 1 71 3 3 1 1 4 3 0
7 67 1 18 1 0 0 1 1 0 0
8 67 1 16 5 4 0 1 1 0 0
9 71 1 20 5 4 0 1 1 0 0
10 67 2 28 4 4 2 1 2 0 0
Table 4. Debugging productivity and quality comparison between traditional bottom-up approach and TapML.
{tblr}
row2,4,6,8=bg=azure9, colspec=Q[c,white] — Q[l] Q[r] Q[c] Q[c] Development Step Platform Cost One-time Effort?